# Training Gradient Boosted Decision Trees on Tabular Data Containing Label Noise for Classification Tasks

### v1

@@ 4 Methodology
This section addresses the different steps taken to prepare the experiment settings with the goal of answering the research questions. First, four methods aimed at the detection of noisy instances are introduced and mathematically defined, followed by two label noise correction methods. Implementation details are also explained in these sections. After that, the datasets used for evaluation are listed and their characteristics briefly summarized, as well as the preprocessing measures taken to prepare the data for modeling. Lastly, the evaluation metrics used in this work and the reasoning behind that selection will be provided.
4.1 Implemented Methods
As mentioned in the “Related Works” section (Section 2), LRT-Correction and AUM Ranking were originally developed for DNNs and were adapted for this work to work on GBDTs. As opposed to Random Forests, which train fully grown trees on subsets of the data and aggregate their predictions using hard or soft voting, GBDTs fit shallow trees to the residual error of the current model, which is usually approximated using the gradient of the current model. Therefore, GBDTs optimize a loss function just as DNNs do, which is why DNN methods can be transferred to GBDTs.
Table 1 summarizes all implemented detection methods and some of their characteristics.
Method Original Domain Data
LRT-Correction DNN Past predicted class probabilities
Training Dynamics Statistics GBDT Class probabilities and predictions
AUM Ranking DNN Past raw margins
Gradients DNN Past gradients
Table 1: An overview of the implemented noise detection methods.
AUM-Ranking (Pleiss et al., 2020) and LRT-Correction (Zheng et al., 2020) use past predictions and calculate a metric related to the consistency of those predictions or a ratio between the predicted and the true class. Gradients isn’t based on a particular method, a combination of the small-loss-trick (Li et al., 2020) and the phenomenon that boosting algorithms tend to assign large weights to noisy instances (Verbaeten and Assche, 2003). These methods were chosen because they all apply to data cleansing, which is within the scope of this work as defined in Section 2.5, and are all well-known in the field of learning with label noise. Training Dynamics Statistics (Ponti et al., 2022) was adapted to an online algorithm, meaning that the metrics are computed at each training iteration instead of once at the end of training and then retraining a model based on these metrics. This change was made to investigate how the metrics change during training and for comparability to the other noise detection methods. Training Dynamics Statistics was chosen because it is one of the few works that investigate learning with label noise using GBDTs and focuses on data cleansing.
Though some methods additionally implement loss correction methods or similar measures, this work will solely focus on noise detection methods in combination with the relabeling and removal of suspicious instances. Also, despite the fact that some of the used datasets are imbalanced, taking class imbalance into account is beyond the scope of this work.
@@ 4.1.1 LRT-Correction
LRT-Correction stands for “likelihood ratio test correction” and is an algorithm that is based on a theorem provided by the authors quantifying how a classifiers prediction based on noisy data correlates to the purity of an instance’s label. Let $p(i|x)$ be the classifier’s predicted probability of class $i\in\{1,\dots,c\}$ and $\hat{y}=\mathrm{argmax}_{i}\ p(i|x)$ the predicted class. Based on the theorem, the ratio $LR(x,\tilde{y})$ between the classifier’s confidence on the noisy label $p(\tilde{y}|x)$ and its own label prediction $p(\hat{y}|x)$ is used to determine whether the instance should be relabeled (Zheng et al., 2020):
$LR(x,\tilde{y})$ $=\frac{p(\tilde{y}|x)}{p(\hat{y}|x)}$ (14)
If this ratio exceeds a certain threshold, i.e. the confidence on the predicted label $\hat{y}$ is much higher than that on the noisy label $\tilde{y}$, the label is flipped to the classifier’s prediction:
$\tilde{y}_{new}=\begin{cases}\hat{y},&\text{if }LR(x,\tilde{y})<\epsilon\\ \tilde{y},&\text{otherwise}\end{cases}.$ (15)
The network is trained using the cross-entropy loss $L_{CE}$. After a few warm-up iterations, an additional retroactive loss term $L_{retro}$ is added to enforce consistency of label predictions in consecutive epochs:
$l(f^{t}(x),\tilde{y},f^{t-1}(x))=L_{\text{retro}}(f^{t}(x),f^{t-1}(x))+L_{CE}(f^{t}(x),\tilde{y})$ (16)
$=\sum_{i=1}^{c}p^{t-1}(i|x)\log p(i|x)+\sum_{i=1}^{c}q_{\tilde{y}}(i)\log p(i|x),$
where $q_{\tilde{y}}(i)=[\tilde{y}=i]$ and where $p^{t-1}(i|x)$ is the probability of class $i$ as predicted by the previous model. The square bracket in $[\tilde{y}=i]$ is the Iverson bracket, which returns 1 if the condition is met and 0 otherwise. In this paper, the likelihood ratio $LR(x,\tilde{y})$ is used to identify noisy labels and either combined with LRT’s label flipping scheme as described in equation (15) or removing them. More on combining methods is described in Section 4.1.5.
@@ 4.1.2 Training Dynamics Statistics
Training dynamics describe the model’s behavior as its training progresses. In data cartography, statistics calculated based on training dynamics, i.e. the predicted probability $p(i|x)$ of class $i$ and predictions $\hat{y}$ across different estimators, are used to categorize data instances into “easy-to-learn”, “hard-to-learn” and “ambiguous”. Confidence $\mu(x)$ measures the average probability for the true label $\tilde{y}$ across all iterations. Variability $\sigma(x)$ measures the fluctuation of an instance’s true class probabilities during training and correctness $\gamma(x)$ defines the percentage of iterations in which the model correctly labels $x$ (Swayamdipta et al., 2020):
@@
with $T$ as the total number of trees (estimators).
Ponti et al. (2022) applied the idea of dataset cartography to GBDTs. They used the product of confidence and correctness to remove problematic instances either by threshold or an instance re-weighting scheme. In this work, we build upon the thresholding method, but in addition to removing problematic instances, also attempt to relabel them. The method will be denoted as ConfCorr in the Experiment section (Section 5).
4.1.3 AUM Ranking
“Area Under the Margin” (AUM) Ranking uses the margins at each epoch to differentiate easy, hard and mislabeled instances. The margin $M$ at epoch $t$ measures how much larger the assigned logit $z^{t}_{\tilde{y}}(x)$ is than the largest logit of another class and $\operatorname{AUM}$ is the mean margin of an instance over all epochs (Pleiss et al., 2020):
A positive margin indicates that a neural network is confident in its prediction. Conversely, a negative margin signifies a lack of confidence or even an incorrect prediction. In their work, Pleiss et al. (2020) additionally move threshold instances to an extra class so that the algorithm can learn to isolate mislabeled instances better from misclassified but correctly labeled instances. This work focuses on using the AUM metric as a means of identifying noisy instances.
@@ 4.1.4 Gradients
This method is based both on the “small-loss-trick” from deep learning as well as the growing weights of noisy instances in boosting algorithms (Verbaeten and Assche, 2003). With the small-loss-trick, small-loss samples from the training set are assumed to be clean and selected to update the network robustly (Han et al., 2018; Jiang et al., 2018; Li et al., 2020). Gradient boosting doesn’t assign weights to instances as standard boosting algorithms do, but the gradients with respect to each instance can be used instead.
An early experiment on the Covertype dataset confirms this: Figure 2 shows the maximum absolute gradient of clean and noisy instances per epoch. After just ten epochs the gradients of noisy instances are on average 1.5 times higher than the gradients of clean instances.
Figure 2: Maximum absolute gradients of noisy and clean instances per epoch (Covertype). The gradient is significantly larger for noisy instances.
In this work, the approach of Li et al. (2020) is followed, where they fit a two-component Gaussian Mixture Model (GMM) to the per-instance loss distribution over the last few epochs. The component with smaller mean is the clean probability. Since the selected objective function (see Section 5.1.2) returns the gradients with respect to each class per epoch, the largest absolute gradient of each instance was used to represent it.
4.1.5 Noise Correction
To limit the scope of this work, the correction methods were limited to removal and relabeling of noisy instances. As mentioned in the related works section (Section 2), other methods could include instance reweighting and loss correction.
All of the above methods were used for noisy instance detection and combined with either removal or relabeling for noisy instance correction, which amount to eight possible combinations. The percentage number of instances that may be removed was set to 80% so that the correction methods don’t accidentally remove almost all of the data, making further learning impossible. In the future, this percentage could be selected dynamically based on the estimated noise rate. For relabeling, the class with the highest probability averaged over the size of the history window (see Section 5.1.3) was chosen as the new label. Also, each instance could only be relabeled once to avoid error accumulation.
@@ 4.2 Data Sets
The model’s performance will be evaluated on four well-known public datasets, namely the UCI Covertype  [^1] https://archive.ics.uci.edu/dataset/31/covertype, accessed 30th July 2024, Dry Bean  [^2] https://archive.ics.uci.edu/dataset/602/dry+bean+dataset, accessed 30th July 2024, Adult  [^3] https://archive.ics.uci.edu/dataset/2/adult, accessed 30th July 2024 and Breast Cancer  [^4] https://archive.ics.uci.edu/dataset/17/breast+cancer+wisconsin+diagnostic, accessed 30th July 2024 dataset. They were chosen for comparability to other studies investigating the use of GBDTs on data containing label noise and because they are suitable for classification tasks.
The Covertype dataset contains cartographic descriptions of forest covers like elevation and soil type and is used to predict the type of forest cover (e.g. “pine”, “spruce” etc.). The labels were obtained from geospatial data from the US Forest Services. The features in the Dry Bean dataset were derived from high-resolution images of dry beans and are used to predict the type of the dry bean. The dry beans used for the images were obtained from certified seed producers (Koklu and Özkan, 2020), meaning that the data is professionally labeled. The Adult dataset is “a set of reasonably clean records” from the 1994 Census database and uses demographic information to predict whether a person’s income exceeds \$50K per year. Similar to the Dry Bean dataset, the Breast Cancer dataset contains features computed from images. The images display the fine needle aspirate (a procedure for obtaining cells from an internal organ or tumors) of breast mass and the features describe the characteristics of the cell nuclei present in the image. This information can be used to predict whether the samples contain benign or malignant cells, i.e. for cancer diagnosis.
Every dataset exhibits a certain level of class imbalance. For the Adult dataset, the class distribution is 76.07% to 23.93%. In the Breast Cancer dataset, the classes are divided into 37.26% and 62.74%. The Dry Bean dataset has its largest class representing 26.05% of the instances, while its smallest class accounts for 3.84%. Among these, the Covertype dataset is the most skewed, with the predominant class constituting 48.76% and the least common class comprising just 0.47% of the total data. The aforementioned proportions are visualized in Figure 3. The smallest class of the Covertype dataset consists of such few instances that its bar is barely visible in the figure.
Figure 3: The percentage of instances in the largest and smallest class per dataset. All datasets possess some degree of class imbalance.
The size of the datasets is also very diverse. The Covertype is the largest dataset, consisting of more than 500 thousand instances, whereas the Breast Cancer dataset merely contains 569. The specific dataset sizes can be viewed in Table 2 and are visualized in Figure 4.
Figure 4: The number of instances in each dataset. The size of the datasets varies a lot.
The labels of all public datasets are assumed to be almost clean, as they were either self-reported (e.g., the Adult dataset), based on other reliable datasets (e.g., the Covertype dataset), or labeled by professionals (e.g., the Dry Bean and Breast Cancer datasets). Characteristics of the datasets are summarized in Table 2.
@@
Table 2: Characteristics of the datasets used in this work.
All datasets were preprocessed by imputing missing data with the median of the feature if it was numeric and with mode otherwise. Numeric attributes were standardized to have a mean of zero and a standard deviation of one and categorical attributes were one-hot encoded. In some cases, features that leaked information about the target variable were discarded to prevent positively skewed results.
4.3 Metrics
In most papers, classification performance is evaluated using accuracy on the test set at various noise rates (Li et al., 2020; Xiang et al., 2020; Zheng et al., 2020). Here, accuracy on the train set is evaluated using the noisy labels and accuracy on the test set, since it isn’t polluted by noise, on the true labels only. Because this paper builds upon Ponti et al.’s (2022) work, precision, recall and the f1-score will also be included for greater comparability of the classification performances. Similarly, the noise detection performance will be evaluated using accuracy, the false negative rate (FNR) and the false positive rate (FPR).
5 Experiments
This section details the different experiment settings set up for answering the research questions and presents the information obtained in them. By providing clarity on the experimental procedures, this section ensures that the research is transparent and can be replicated by others. First, the experimental conditions are specified. This entails an explanation of the noise injection process, information about which noise types are used and the amount of noise that was injected into the data. The gradient-boosting library that was used as well as the model and other training parameters like early stopping are specified as well. The noise detection and correction methods used in the experiments are listed after that, which have already been introduced in Section 4.
The first section with experimental results covers the first set of research questions regarding the effects of label noise on GBDTs with no account for label noise. The experiments show that GBDTs are robust to label noise in the early epochs of training, especially to symmetric noise. They mostly predict the ground-truth true label even if it was altered due to label noise. Symmetric noise is generally harder to learn than pair noise but subsequent performance on the test set is higher.
The next section covers the second set of research questions, which are about the performance of the noise detection and correction methods and how they are influenced by the amount of noise present in the data. The experiments show that AUM and LRT perform best in terms of noise detection accuracy, closely followed by ConfCorr and Gradients. The percentage of instances marked as noisy by AUM and LRT can even be used to get a sense of the amount of noise present in the data. Combining these methods with either relabeling or removing instances marked as noisy results in a lower classification loss, but classification accuracy worsens. It could be that relabeling introduces more errors in the data and removing excludes hard to learn but informative instances from the training set.
To supplement the visual comparison of all applied methods, the “Final Results” section (Section 5.4) presents the results of different experimental settings in tabular form and compares them with other studies. AUM, LRT, and Gradients are demonstrated to achieve state-of-the-art noise detection accuracy on the Adult dataset. Additionally, ConfCorr and Gradients, when combined with noisy candidate removal, attain the highest recall score on the Breast Cancer dataset compared to other studies. ConfCorr, when combined with relabeling, achieved the highest precision score on the Adult dataset compared to other works. The most effective combination of noise detection and correction methods varies depending on the dataset, noise rate, and noise type. However, in terms of noise detection and classifier performance, the implemented methods were generally outperformed by other state-of-the-art work.
5.1 Experiment Conditions
5.1.1 Noise Injection
@@
Assuming that all datasets are sufficiently clean, a noise transition matrix for pair and symmetric noise was applied to the training labels as defined in Section 3.2. The methods were evaluated on seven noise rates, $\tau\in\{0.0,0.1,0.2,0.3,0.4,0.5,0.6\}$, for each type of noise. In the case of pair noise, noise rates exceeding 50% make misclassification more likely than the correct label. However, these higher rates were still included to assess their impact on the model’s performance. In the final comparison (Section 5.4), only noise rates ranging from 10% to 40% were considered.
5.1.2 Model and Hyperparameters
@@
For the classification task, a GBDT model was trained using the XGBoost library. The default parameters proved to be the most effective: maximum tree depth was set at 6, learning rate at 0.3, and the tree method was auto. Because the low-level boosting interface was used, early stopping had to be implemented manually with a minimum required loss improvement of 0.5 and a patience of 10 epochs. The model optimized the cross-entropy objective function and had a warm-up duration of 15 epochs. In deep learning, warm-up refers to the initial part of training where hyperparameters, especially the learning rate, are gradually increased from a lower value to the target value. This approach is employed to avoid instability or divergence of the training process, ensuring a more stable and effective convergence of the model. In the context of GBDTs, the term “warm-up” will be used to describe training the model for a few iterations to ensure that noise detection methods have enough past predictions to calculate their metrics from.
@@
XGBoost’s softprob and logistic objective functions were used for multiclass and binary classification, respectively. softprob is essentially the cross-entropy loss function $H$, which is based on the softmax function $P_{\sigma}$ and returns the probability of each class per instance (Géron, 2019):
$P_{\sigma}(y=i)$ $=\frac{e^{z_{i}}}{\sum_{j=1}^{c}e^{z_{j}}},$ (22)
$H_{\sigma}(y)$ $=-\sum_{i=1}^{N}\sum_{k=1}^{c}[y_{i}=k]\cdot\log(P_{\sigma}(y_{i}=k)),$ (23)
where $z_{i}$ is the logit for class $i$. Logistic uses logistic regression as the objective. It is defined using the logistic function $P_{\kappa}$ to map the logits to a probability value between zero and one and again combined with cross-entropy $H_{\kappa}$ as the loss function:
$P_{\kappa}(y=1)=\frac{1}{1+e^{-z}},$ (24)
$H_{\kappa}(y)=\sum_{i=1}^{N}\left[y_{i}\log(P_{\kappa}(y_{i}=1))+(1-y_{i})\log(1-P_{\kappa}(y_{i}=1))\right].$ (25)
5.1.3 Algorithms
The detection methods LRT, ConfCorr, AUM and Gradients combined with the correction methods remove and relabel from Section 4 are compared with a baseline algorithm (denoted by none), which is a GBDT model trained without any noise treatment methods.
@@ 5.1.3 Algorithms
As in the original implementation, ConfCorr calculates its metrics incrementally across all epochs. The remaining noise detection methods use a history of five epochs to calculate their metrics from.
As outlined in Section 4.1.5, the removal of noisy labels was capped at 80% of the training data. If a detection algorithm identified more instances, the excess instances were retained. Each instance was allowed to be relabeled only once during relabeling to prevent the accumulation of errors.

### v2

@@ 4 Methodology
This section presents the noise detection and correction methods implemented and adapted for this work.
4.1 Noise Detection Methods
Four methods were implemented for detecting noisy instances: LRT-Correction, AUM Ranking, Training Dynamics Statistics, and Gradients. LRT-Correction and AUM Ranking were adapted from deep neural networks (DNNs), while Training Dynamics Statistics is based on prior GBDT research. The Gradients method, a novel approach here, is inspired by weight clipping in Adaboost and the small-loss trick.
@@ 4.1.1 LRT-Correction
LRT-Correction (Likelihood Ratio Test Correction) uses a likelihood ratio to evaluate the purity of an instance’s label. Let $p(i|x)$ denote the classifier’s predicted probability for class $i$, and $\hat{y}=\arg\max_{i}p(i|x)$ represent the predicted class. The likelihood ratio $LR(x,\tilde{y})=\frac{p(\tilde{y}|x)}{p(\hat{y}|x)}$ is calculated between the classifier’s confidence in the noisy label $\tilde{y}$ and its prediction $\hat{y}$ (Zheng et al., 2020):
$\tilde{y}_{new}=\begin{cases}\hat{y},&\text{if }LR(x,\tilde{y})<\epsilon,\\ \tilde{y},&\text{otherwise}.\end{cases}$ (1)
The original approach also includes an additional retroactive loss term to improve label consistency across epochs, although this term was not implemented in this study. Here, the likelihood ratio $LR(x,\tilde{y})$ is used to identify noisy labels, combined with either relabeling or removal of these instances (see Section 4.2).
4.1.2 AUM Ranking
”Area Under the Margin” (AUM) Ranking uses margins at each epoch to classify instances as easy, hard, or mislabeled. The margin $M$ at epoch $t$ measures the difference between the assigned logit $z^{t}_{\tilde{y}}(x)$ and the highest logit of another class, with $\operatorname{AUM}$ representing the average margin over all epochs (Pleiss et al., 2020):
A positive margin indicates confidence in prediction, while a negative margin suggests uncertainty or mislabeling. Although Pleiss et al. (2020) also categorize threshold instances into a separate class to better isolate mislabeled data, this work focuses on the AUM metric solely for identifying noisy instances.
@@ 4.1.3 Training Dynamics Statistics
Training dynamics capture a model’s behavior during training and are used in data cartography to classify instances as ”easy-to-learn,” ”hard-to-learn,” or ”ambiguous.” This approach uses the predicted probability $p(i|x)$ and predictions $\hat{y}$ across estimators $t$. Confidence $\mu(x)$ is the average probability for the true label $\tilde{y}$, variability $\sigma(x)$ measures fluctuation in true class probabilities, and correctness $\gamma(x)$ represents the percentage of correct classifications (Swayamdipta et al., 2020):
@@
where $T$ is the number of trees.
Ponti et al. (2022) applied dataset cartography to GBDTs, using confidence and correctness to filter problematic instances through thresholding or re-weighting. Here, we extend the thresholding method to also relabel identified instances, denoted as ConfCorr in Section 5.
@@ 4.1.4 Gradients
This method combines the ”small-loss trick” from deep learning with the growing gradients of noisy instances observed in boosting algorithms (Verbaeten and Assche, 2003). The small-loss trick assumes that low-loss samples are likely clean, updating the network with these instances (Han et al., 2018; Jiang et al., 2018; Li et al., 2020). Although gradient boosting does not assign instance weights, the per-instance gradients serve a similar purpose.
An experiment on the Covertype dataset shows that noisy instances display larger gradients than clean instances, with noisy gradients averaging 1.5 times higher by epoch ten (see Figure 2).
Figure 2: Maximum absolute gradients of noisy and clean instances per epoch (Covertype). Noisy instances exhibit significantly larger gradients.
Following Li et al. (2020), a two-component Gaussian Mixture Model (GMM) is fitted to the loss distribution over the final epochs, with the component of smaller mean indicating clean probabilities. The largest absolute gradient per instance represents its noisiness in this model.
4.2 Noise Correction
This work focuses on two noise correction methods: removal and relabeling. While other methods like instance reweighting and loss correction exist, they are beyond this study’s scope.
Each noise detection method was paired with either removal or relabeling, yielding eight combinations. For relabeling, instances were reassigned to the class with the highest probability averaged over a history window and could only be reassigned once. To preserve sufficient data for training, removal was limited to at most 80% of the entire dataset.
5 Experiments
This section outlines the datasets, noise injection process, evaluation metrics, model configuration, and experimental settings used in this study.
@@ 5.1 Datasets
This study evaluates model performance on four public datasets frequently used in label noise and GBDT research: Covertype  [^1] https://archive.ics.uci.edu/dataset/31/covertype, accessed 30th July 2024, Dry Bean  [^2] https://archive.ics.uci.edu/dataset/602/dry+bean+dataset, accessed 30th July 2024, Adult  [^3] https://archive.ics.uci.edu/dataset/2/adult, accessed 30th July 2024 and Breast Cancer  [^4] https://archive.ics.uci.edu/dataset/17/breast+cancer+wisconsin+diagnostic, accessed 30th July 2024. These datasets represent diverse classification tasks and data types, as summarized in Table 1. Each dataset varies significantly in size and degree of class imbalance, with Covertype being the largest and most imbalanced.
@@
Table 1: Characteristics of the datasets used in this study.
The labels are assumed to be nearly clean, given their sources: Covertype and Adult rely on reliable datasets, while Dry Bean and Breast Cancer are professionally labeled. Preprocessing included imputing missing values, standardizing numeric attributes, and one-hot encoding categorical features.
5.2 Noise Injection
@@
Assuming that all datasets are sufficiently clean, a noise transition matrix for pair and symmetric noise was applied to the training labels as defined in Section 3. The methods were evaluated on seven noise rates, $\tau\in\{0.0,0.1,0.2,0.3,0.4,0.5,0.6\}$, for each type of noise. In the case of pair noise, noise rates exceeding 50% make misclassification more likely than the correct label. However, these higher rates were still included to assess their impact on the model’s performance. In the final comparison (Section 6.3), only noise rates ranging from 10% to 40% were considered.
5.3 Evaluation Metrics
Classification performance is evaluated using test set accuracy, precision, recall, and F1-score, with test accuracy based on true labels. Training accuracy, measured with noisy labels, is shown only in the experimental figures, not in the final comparison to other works. For noise detection, accuracy is reported because it is a widely used metric across most papers, ensuring comparability with prior research.
5.4 Model and Hyperparameters
For the classification task, a GBDT model was trained using the XGBoost library. The default parameters proved to be the most effective: maximum tree depth was set at 6, learning rate at 0.3, and the tree method was auto.
@@ 5.4 Model and Hyperparameters
Because the low-level boosting interface was used, early stopping had to be implemented manually with a minimum required loss improvement of 0.5 and a patience of 10 epochs. The model optimized the cross-entropy objective function and had a warm-up duration of 15 epochs. In deep learning, warm-up refers to the initial part of training where hyperparameters, especially the learning rate, are gradually increased from a lower value to the target value. This approach is employed to avoid instability or divergence of the training process, ensuring a more stable and effective convergence of the model. In the context of GBDTs, the term “warm-up” is be used to describe training the model for a few iterations to ensure that noise detection methods have enough past predictions to calculate their metrics from.
@@
XGBoost’s softprob and logistic objective functions were used for multiclass and binary classification, respectively.
@@
Consistent with the original ConfCorr implementation, metrics are calculated incrementally across all epochs. The remaining noise detection methods calculate metrics over a history of five epochs.
5.5 Experimental Settings
The experimental setup is divided into three stages. In the first stage, the impact of label noise on GBDTs is assessed by training the models on datasets with noisy labels and evaluating them on clean datasets.
In the second part of the experiments, the noise detection and correction methods are applied and compared to each other. Each detection method is paired with each correction method and evaluated against a baseline model (denoted by none), which is a GBDT trained without any noise-handling measures and serves as a control to gauge the effects of noise detection and correction by comparing them to an uncorrected model.
Finally, in the third stage, the performance of the implemented methods are compared to the work of Ponti et al. (2022).
6 Results
The following sections present the results of experiments on label noise in GBDTs. The impact of label noise without correction is first examined, followed by a comparison of the implemented noise detection and correction techniques. Finally, these methods are compared to a state-of-the-art approach.